The Arrhenius equation is k = Ae–Ea/RT. The slope of a plot of ln k vs. 1/T is equal to

A) -k

B) k

C) Ea

D) -Ea/R

E) A
Answer #1

Taking the log of each side

In k =In(Ae-Ea/Rt) =In A+ In(e-Ea/Rt) =In A + (-Ea/Rt)

In k=--Ea/Rt + In A

Which can be vuewvi as

In k-(-Ea/R) (1/T) + In A

This is in the form y=mx+b where x=1/T and y= In k

So the slope of the straight line is

-Ea/R answer D